A 11.56 g mixture of sugar (C_12H_22O_M) and table
Answer #1

delta T = i*Kf*m

Kf of water = -1.86

i = 3

NaCl = Na+ + Cl-

and sugar is a non electrolyte = 1

i = 2+1 = 3

-2.57 = 3* [-1.86]*m

m = 0.46057

Molality = moles of solute / Mass of solvent in Kg

0.46057 = moles of solute / 0.203

Moles of solute = 0.0935

Moles of solute = moles of Sugar + Moles of NaCl

Moles of sugar = Mass /Molecular weight = 0.0338

Moles of NaCl= 0.0935 - 0.0338 = 0.0597

mass of NaCl = 0.0597*58.44 = 3.488 gms

% percent of sugar = 11.56 *100 / [11.56+3.488+203] = 5.3 %